III. Suppose a substance has been prepared that is composed of carbon, hydrogen, and nitrogen. When 0.1156 gram of this compound is reacted with oxygen, 0.1638 gram of carbon dioxide (CO2) and 0.1676 gram of water (H2O) are collected. If all the carbon in the compound is converted to CO2




1. Determine the mass of Carbon and hydrogen in the sample




1
Expert's answer
2021-11-17T11:41:02-0500

Moles of C=0.1638/44=3.722×10-3

Mole of H =0.1676/18×2=0.01862

Mass of C= 0.0446g